OpenStudy (anonymous):

what is the equivalent representation of 3^-4 ?

OpenStudy (luigi0210):

\[\LARGE a^{-x}=\frac{1}{a^x}\]

OpenStudy (anonymous):

a. 1/81 b. 1/9 c. 1/27 d. 1/12

OpenStudy (anonymous):

those r the answer choices

OpenStudy (luigi0210):

I'm not giving you an answer, you figure it out, I'm just here to guide.. so: \[\LARGE \frac{1}{3^4}=\frac{1}{(3*3)*(3*3)}=\frac{1}{9*9}=..?\]